# On Closed-Form Expressions for the Fisher–Rao Distance

### v2

@@ 1 Introduction
Information geometry [1, 16, 49] uses the tools of differential geometry to study spaces of probability distributions by regarding them as differential manifolds, called statistical manifolds. When these distributions are parametric, a structure of interest is the Fisher metric, a Riemannian metric induced by the Fisher information matrix. This is essentially the unique Riemannian metric on statistical manifolds that is invariant by sufficient statistics [6, 42], making it a natural choice to study the geometry of these manifolds. Moreover, this structure allows one to define the Fisher–Rao distance between two probability distributions on the statistical manifold as the geodesic distance between them, i.e., the length of the minimising path, according to the Fisher metric.
@@
However, finding closed-form expressions for the Fisher–Rao distance of arbitrary distributions, however, is not a trivial task—as a matter of fact, more generally, finding geodesics in an arbitrary manifold is a hard problem in differential geometry. Having that in mind, in this work, we collect examples of statistical models for which closed-form expressions for the Fisher–Rao distance are available. Most of them have been published over the last decades, in different places, and we aim to present these results in a unified and accessible language, hoping to bring them to a broader audience. In curating this collection, we also add a few contributions, such as: illustrating the relation between the manifold of negative multinomial distributions and the hyperbolic model; including a few new examples (Rayleigh, Erlang, Laplace, generalised Gaussian, power function, inverse Wishart), to the best of the authors’ knowledge; and writing more examples in the standard form of univariate elliptical distributions (Laplace, generalised Gaussian, logistic). Finally, we note that numerical methods have been proposed to compute the Fisher–Rao distance when no closed-form expression is available, as in [4, 30, 38, 39, 51, 62, 63], but those techniques are beyond the scope of the present work.
We review preliminaries of information geometry in Section 2 and of hyperbolic geometry in Section 3. We then collect closed-form expressions for discrete distributions in Section 4, and for continuous distributions in Section 5. Product distributions are discussed in Section 6, and Section 7 concludes the paper.
@@
We denote the sets =$\mathbb{N}=\{0,1,2,\dots\}$, *=$\mathbb{N}^{*}=\{1,2,3,\dots\}$, $\mathbb{R}_{+}=\left[0,\infty\right[$, and +*$\mathbb{R}_{+}^{*}=\left]0,\infty\right[$. $\mathbbm{1}_{A}(x)$ is the indicator function, that takes value $1$ if $x\in A$, and 00 otherwise. $\delta_{ij}\coloneqq\mathbbm{1}_{\{j\}}(i)$ denotes the Kronecker delta. We denote $\dot{x}(t)\coloneqq\frac{\mathrm{d}}{\mathrm{d}t}x(t)$. $P_{n}(\mathbb{R})$ denotes the cone of $n\times n$ real symmetric definite-positive matrices.
2 Information Geometry Preliminaries
@@
$\mathcal{S}\coloneqq\left\{p_{\xi}=p(x;\xi)\;\middle|\;\xi=(\xi^{1},\dots,\xi^{n})\in\Xi\subseteq\mathbb{R}^{n}\right\},$ (2.1)
is a family of probability distributions $p_{\xi}$ parametrised by $n$-dimensional vectors $\xi=\left(\xi^{1},\dots,\xi^{n}\right)$ such that the mapping $\xi\mapsto p_{\xi}$ is injective and $\Xi$ is an open set of $\mathbb{R}^{n}$. We consider statistical models in which the support of $p_{\xi}$ does not depend on $\xi$, and we take $\mathcal{X}=\operatorname{supp}p_{\xi}$, unless otherwise stated. Note that $\mathcal{S}$ is contained in the infinite-dimensional space $\mathcal{P}(\mathcal{X})\coloneqq\left\{p\in L^{1}(\mu)\;\middle|\;p>0,\ \int_{\mathcal{X}}p~{}\mathrm{d}\mu=1\right\}$ of positive, $\mu$-integrable functions of unit total measure.
@@
To introduce a differentiable structure in $\mathcal{S}$, we consider the following assumptions: 1) the parametrisation $\varphi\colon\Xi\to\mathcal{P}(\mathcal{X}),\ \varphi(\xi)=p_{\xi}$ is a homeomorphism on its image; 2) denoting $\partial_{i}\coloneqq\frac{\partial}{\partial\xi^{i}}$, the functions $\left\{\partial_{1}p_{\xi},\dots,\partial_{n}p_{\xi}\right\}$ are linearly independent; 3) the mapping $\xi\mapsto p_{\xi}(x)$ is smooth, for all $x\in\mathcal{X}$; 4) the partial derivatives $\partial_{i}p_{\xi}(x)$ commute with the integrals. Moreover, by considering diffeomorphic parametrisations as equivalent, $\mathcal{S}$ becomes a differentiable manifold, called statistical manifold. Note that the parametrisation $\xi\mapsto p_{\xi}$ is a global coordinate system for this manifold (see Figure 1).
@@
Alternatively, the Fisher matrix can be written as the negative expectation of the Hessian of the log-likelihood function:
Proposition 2.1 ([16, Prop. 1.6.3]).
@@
The next two results state important properties of the Fisher metric.
Proposition 2.2 ([16, Thm. 1.6.4]).
@@
Proposition 2.3 ([16, Thm. 1.6.5]).
Proof.
These properties make the Fisher metric a natural choice of Riemannian metric for statistical manifolds. In fact, this is the unique metric (up to a multiplicative constant) in statistical manifolds that is invariant under sufficient statistics [6, 42].
Furthermore, the Fisher metric induces a notion of distance between two distributions in the same statistical manifold, called Fisher–Rao distance, and given by the geodesic distance between these points. Specifically, consider a curve $\xi\colon\left[0,1\right]\to\Xi$ in the parameter space and its image $\gamma\colon\left[0,1\right]\to\mathcal{S}$ by the parametrisation $\varphi$, i.e., $\gamma(t)=(\varphi\circ\xi)(t)$, for $t\in\left[0,1\right]$. Since $\dot{\gamma}(t)=\mathrm{d}\varphi_{\xi(t)}\big{(}\dot{\xi}(t)\big{)}$, in the Fisher geometry, the length of $\gamma$ can be computed as
@@
Then, given two distributions $p_{\xi_{1}}$ and $p_{\xi_{2}}$ in $\mathcal{S}$, the Fisher–Rao distance between them111 We shall also refer to the Fisher–Rao distance between two parametric distributions as the distance between their respective parameters. is the infimum of the length of piecewise differentiable curves $\gamma$ joining these two points:
@@
$\ddot{\xi}^{k}(t)+\sum_{i=1}^{n}\sum_{j=1}^{n}\Gamma_{ij}^{k}\dot{\xi}^{i}(t)\dot{\xi}^{j}(t)=0,\quad k=1,\dots,n,$ (2.9)
where $\Gamma_{i,j}^{k}$ are the Christoffel symbols of the second kind, which can be obtained by the equation
$\sum_{k=1}^{n}g_{\ell k}\Gamma_{ij}^{k}=\frac{1}{2}\left(\frac{\partial}{\partial\xi^{i}}g_{j\ell}+\frac{\partial}{\partial\xi^{j}}g_{\ell i}-\frac{\partial}{\partial\xi^{\ell}}g_{ij}\right).$
The Hopf–Rinow theorem (e.g., [34, Thm. 6.4.6]) provides a sufficient condition for the minimum length, as in (2.8), to be realised by a geodesic: if $(\mathcal{S},d_{\mathrm{FR}})$ is connected and complete as a metric space, then any two points $p,q$ in the manifold $\mathcal{S}$ can be joined by a minimising curve which is a geodesic, that is, a curve whose length is equal to the Fisher–Rao distance $d_{\mathrm{FR}}(p,q)$. This condition is satisfied in all statistical manifolds considered in this paper.
@@
The Fisher–Rao distance is related to the Kullback-Leibler divergence [16, Thm. 4.4.5]:
@@
For higher-dimensional manifolds, the techniques to find the geodesics and the Fisher–Rao distance consist in directly solving the geodesics differential equations, or in doing an analogy with other well-known geometry (e.g., spherical, hyperbolic), as we shall see in the next sections.
@@ 3 Hyperbolic Geometry Results
We recall in this section some classical results from hyperbolic geometry [9, 20, 61], since many of the statistical manifolds studied in this work are related to that geometry. These will be extensively used particularly in § 4.7 and § 5.45.9. If desired, the reader may skip this section for now, and return to it when reading those subsections, to get the details of the derivations.
We start with the hyperbolic geometry in dimension two, to be used in approaching the statistical manifolds in § 5.45.9, analogously to what is done in [22]. In this case, we consider the Poincaré half-plane $\mathcal{H}^{2}\coloneqq\left\{(x,y)\in\mathbb{R}^{2}\;\middle|\;y>0\right\}$ as a model for hyperbolic geometry, with the metric given in matrix form by
@@
with $a,b$ positive constants, then the geodesic distances between two points in $\mathcal{M}$ is given by
@@
On the other hand, consider the diffeomorphism $\psi\colon\varphi(x,y)\in\mathcal{M}\mapsto(\sqrt{a}x,\sqrt{b}y)\in\mathcal{H}^{2}$, and the image curve $\beta(t)\coloneqq(\psi\circ\alpha)(t)=(\sqrt{a}x(t),\sqrt{b}y(t))$. Applying $G_{\mathcal{H}^{2}}(\beta(t))$ to $\dot{\beta}(t)=\mathrm{d}\psi_{\alpha(t)}(\dot{\alpha}(t))$ in $\mathcal{H}^{2}$ gives
@@
Thus, $\left\langle\dot{\alpha}(t),\dot{\alpha}(t)\right\rangle_{G_{\mathcal{M}}(\alpha(t))}=b\left\langle\dot{\beta}(t),\dot{\beta}(t)\right\rangle_{G_{\mathcal{H}^{2}}(\beta(t))}$, that is, $\left\|\dot{\alpha}(t)\right\|_{G_{\mathcal{M}}(\alpha(t))}=\sqrt{b}\left\|\dot{\beta}(t)\right\|_{G_{\mathcal{H}^{2}}(\beta(t))}$. This implies that a curve $\alpha(t)$ connecting two points in $\mathcal{M}$ is a geodesic (minimises the length), if only if, its image $\beta(t)$ is a geodesic in $\mathcal{H}^{2}$. Taking (2.7) and (2.8) into account concludes the proof. ∎
@@
that is, $\left\langle(u_{1},\dots,u_{n}),(v_{1},\dots,v_{n})\right\rangle_{\mathcal{G}_{\mathcal{H}^{n}}(x_{1},\dots,x_{n})}=\frac{1}{x_{n}^{2}}\sum_{i=1}^{n}u_{i}v_{i}$. This manifold has constant negative curvature. The geodesics in this manifold are vertical half-lines and vertical half-circles centred at the hyperplane $x_{n}=0$, and the geodesic distance between two points in $\mathcal{H}^{n}$ is given by
@@
Restricted to points in the $n$-dimensional unit half-sphere in $\mathcal{H}^{n+1}$, that is, $S_{1}^{n}\coloneqq\left\{(x_{1},\dots,x_{n+1})\in\mathcal{H}^{n+1}\;\middle|\;\sum_{i=1}^{n+1}x_{i}^{2}=1\right\}$, the expression for the distance becomes
$d_{\mathcal{H}^{n+1}}\big{(}(x_{1},\dots,x_{n+1}),(y_{1},\dots,y_{n+1})\big{)}=\operatorname{arccosh}\left(\frac{1-\sum_{i=1}^{n}x_{i}y_{i}}{x_{n+1}y_{n+1}}\right).$ (3.5)
From (3.4), we immediately see that the distance $d_{\mathcal{H}^{n}}$ is invariant by dilation or contraction, that is, denoting $\boldsymbol{x}\coloneqq(x_{1},\dots,x_{n})\in\mathcal{H}^{n}$ and $\boldsymbol{y}\coloneqq(y_{1},\dots,y_{n})\in\mathcal{H}^{n}$, we have $d_{\mathcal{H}^{n}}(\boldsymbol{x},\boldsymbol{y})=d_{\mathcal{H}^{n}}(\lambda\boldsymbol{x},\lambda\boldsymbol{y})$, for $\lambda\neq 0$. In particular, the expression (3.5) holds for points in the half-sphere $S_{r}^{n}\coloneqq\left\{(x_{1},\dots,x_{n+1})\in\mathcal{H}^{n+1}\;\middle|\;\sum_{i=1}^{n+1}x_{i}^{2}=r\right\}$ of any radius $r$. This will be used in § 4.7.

### v3

@@ 1 Introduction
Information geometry [1, 2, 3] uses the tools of differential geometry to study spaces of probability distributions by regarding them as differential manifolds, called statistical manifolds. When these distributions are parametric, a structure of interest is the Fisher metric, a Riemannian metric induced by the Fisher information matrix. This is essentially the unique Riemannian metric on statistical manifolds that is invariant by sufficient statistics [4, 5], making it a natural choice to study the geometry of these manifolds. Moreover, this structure allows one to define the Fisher–Rao distance between two probability distributions on the same statistical manifold as the geodesic distance between them, i.e., the length of the minimising path, according to the Fisher metric.
@@
However, finding closed-form expressions for the Fisher–Rao distance of arbitrary distributions is not a trivial task—as a matter of fact, more generally, finding geodesics in an arbitrary manifold is a difficult problem in differential geometry. Having that in mind, in this work, we collect examples of statistical models for which closed-form expressions for the Fisher–Rao distance are available. Most of them have been published over the last decades, in different places, and we aim to present these results in a unified and accessible language, hoping to bring them to a broader audience. In curating this collection, we also add a few contributions, such as: illustrating the relation between the manifold of negative multinomial distributions and the hyperbolic model; including a few new examples (Rayleigh, Erlang, Laplace, generalised Gaussian, power function, inverse Wishart), to the best of the authors’ knowledge; and writing more examples in the standard form of univariate elliptical distributions (Laplace, generalised Gaussian, logistic). Finally, we note that numerical methods have been proposed to compute the Fisher–Rao distance when no closed-form expression is available, as in [34, 35, 36, 26, 37, 27, 38], but those techniques are beyond the scope of the present work.
We review preliminaries of information geometry in Section LABEL:sec:preliminaries and of hyperbolic geometry in Section 2. We then collect closed-form expressions for discrete distributions in Section 3, and for continuous distributions in Section 4. Product distributions are discussed in Section 5, and Section 6 concludes the paper.
@@
We denote the sets =$\mathbb{N}=\{0,1,2,\dots\}$, =$\mathbb{N}^{*}=\{1,2,3,\dots\}$, $\mathbb{R}_{+}=\left[0,\infty\right[$, and $\mathbb{R}_{+}^{*}=\left]0,\infty\right[$. $\mathbbm{1}_{A}(x)$ is the indicator function, that takes value $1$ if $x\in A$, and 00 otherwise. $\delta_{ij}\coloneqq\mathbbm{1}_{\{j\}}(i)$ denotes the Kronecker delta. We denote $\dot{x}(t)\coloneqq\frac{\mathrm{d}}{\mathrm{d}t}x(t)$. $P_{n}(\mathbb{R})$ denotes the cone of $n\times n$ real symmetric definite-positive matrices.
@@
$\mathcal{S}\coloneqq\left\{p_{\xi}=p(x;\xi)\;\middle|\;\xi=(\xi^{1},\dots,\xi^{n})\in\Xi\subseteq\mathbb{R}^{n}\right\}$ (1.1)
is a family of probability distributions $p_{\xi}$ parametrised by $n$-dimensional vectors $\xi=\left(\xi^{1},\dots,\xi^{n}\right)$ such that the mapping $\xi\mapsto p_{\xi}$ is injective and $\Xi$ is an open set of $\mathbb{R}^{n}$. We consider statistical models in which the support of $p_{\xi}$ does not depend on $\xi$ and we take $\mathcal{X}=\operatorname{supp}p_{\xi}$, unless otherwise stated. Note that $\mathcal{S}$ is contained in the infinite-dimensional space $\mathcal{P}(\mathcal{X})\coloneqq\left\{p\in L^{1}(\mu)\;\middle|\;p>0,\ \int_{\mathcal{X}}p~{}\mathrm{d}\mu=1\right\}$ of positive, $\mu$-integrable functions of unit total measure.
@@
To introduce a differentiable structure in $\mathcal{S}$, we consider the following assumptions: 1) the parametrisation $\varphi\colon\Xi\to\mathcal{P}(\mathcal{X}),\ \varphi(\xi)=p_{\xi}$ is a homeomorphism on its image; 2) denoting $\partial_{i}\coloneqq\frac{\partial}{\partial\xi^{i}}$, the functions $\left\{\partial_{1}p_{\xi},\dots,\partial_{n}p_{\xi}\right\}$ are linearly independent; 3) the mapping $\xi\mapsto p_{\xi}(x)$ is smooth, for all $x\in\mathcal{X}$; 4) the partial derivatives $\partial_{i}p_{\xi}(x)$ commute with the integrals. Moreover, by considering diffeomorphic parametrisations as equivalent, $\mathcal{S}$ becomes a differentiable manifold, that we may call a statistical manifold111 We follow the nomenclature from [1], but remark that, more generally, statistical manifold refers to a manifold equipped with a Riemannian metric and a $3$-symmetric tensor, from a purely geometric point of view [39] (see also [40, § 4.5]). We will restrict ourselves to the case of parametric statistical models presented above. . Note that the parametrisation $\xi\mapsto p_{\xi}$ is a global coordinate system for this manifold (see Figure 1).
@@
Alternatively, the Fisher matrix can be written as the negative expectation of the Hessian of the log-likelihood function, a result that can make the computation of the Fisher matrix easier in some cases:
Proposition 1.1 ([2, Prop. 1.6.3]).
@@
The following results will help some derivations in the rest of the text. First, we note that, as any Riemannian metric, the Fisher metric is covariant under reparametrisation of the parameter space:
Proposition 1.2 ([2, Thm. 1.6.5]).
Proof.
Second, we state an invariance property specific to the Fisher metric:
Proposition 1.3 ([2, Thm. 1.6.4]).
@@
Furthermore, it can be shown that the Fisher metric is the unique Riemannian metric (up to a multiplicative constant) in statistical manifolds that is invariant under sufficient statistics [4, Thm. 1.2] (see also [5]). This invariance characterisation justifies the choice of this metric to study the geometry of statistical models.
The Fisher metric induces a notion of distance between two distributions in the same statistical manifold, called Fisher–Rao distance, and given by the geodesic distance between these points. Specifically, consider a curve $\xi\colon\left[0,1\right]\to\Xi$ in the parameter space and its image $\gamma\colon\left[0,1\right]\to\mathcal{S}$ by the parametrisation $\varphi$, i.e., $\gamma(t)=(\varphi\circ\xi)(t)$, for $t\in\left[0,1\right]$. Since $\dot{\gamma}(t)=\mathrm{d}\varphi_{\xi(t)}\big{(}\dot{\xi}(t)\big{)}$, in the Fisher geometry, the length of $\gamma$ can be computed as
@@
Then, given two distributions $p_{\xi_{1}}$ and $p_{\xi_{2}}$ in $\mathcal{S}$, the Fisher–Rao distance between them222 We shall also refer to the Fisher–Rao distance between two parametric distributions as the distance between their respective parameters. is the infimum of the length of piecewise differentiable curves $\gamma$ joining these two points:
@@
$\ddot{\xi}^{k}(t)+\sum_{i=1}^{n}\sum_{j=1}^{n}\Gamma_{ij}^{k}\dot{\xi}^{i}(t)\dot{\xi}^{j}(t)=0,\quad k\in\{1,\dots,n\},$ (1.9)
where $\Gamma_{ij}^{k}$ are the Christoffel symbols of the second kind, which can be obtained from the equations
$\sum_{k=1}^{n}g_{\ell k}\Gamma_{ij}^{k}=\frac{1}{2}\left(\frac{\partial}{\partial\xi^{i}}g_{j\ell}+\frac{\partial}{\partial\xi^{j}}g_{\ell i}-\frac{\partial}{\partial\xi^{\ell}}g_{ij}\right),\quad i,j,\ell\in\{1,\dots,n\}$
The Hopf–Rinow theorem (e.g., [41, Thm. 6.4.6]) provides a sufficient condition for the minimum length, as in (1.8), to be realised by a geodesic: if $(\mathcal{S},d_{\mathrm{FR}})$ is connected and complete as a metric space, then any two points $p,q$ in the manifold $\mathcal{S}$ can be joined by a minimising curve which is a geodesic, that is, a curve whose length is equal to the Fisher–Rao distance $d_{\mathrm{FR}}(p,q)$. This condition is satisfied in all statistical manifolds considered in this paper.
@@
The Fisher–Rao distance is related to the Kullback-Leibler divergence [2, Thm. 4.4.5]:
@@
For higher-dimensional manifolds, the techniques to find the geodesics and the Fisher–Rao distance consist in directly solving the geodesic differential equations, or in doing an analogy with some well-known geometry (e.g., spherical, hyperbolic), as we shall see in the next sections.
@@ 2 Hyperbolic Geometry Results
We recall in this section some classical results from hyperbolic geometry [42, 43, 44], since many of the statistical manifolds studied in this work are related to that geometry. These will be extensively used particularly in § 3.7 and § 4.44.9. If desired, the reader may skip this section for now, and return to it when reading those subsections to get the details of the derivations.
We start with the hyperbolic geometry in dimension two, to be used in the approach of the statistical manifolds in § 4.44.9, analogously to what is done in [45]. In this case, we consider the Poincaré half-plane $\mathcal{H}^{2}\coloneqq\left\{(x,y)\in\mathbb{R}^{2}\;\middle|\;y>0\right\}$ as a model for hyperbolic geometry, with the metric given in matrix form by
@@
with $a,b$ positive constants, then the geodesic distance between two points in $\mathcal{M}$ is
@@
On the other hand, consider the diffeomorphism $\psi\colon\mathcal{M}\to\mathcal{H}^{2}$, $\varphi(x,y)\mapsto(\sqrt{a}x,\sqrt{b}y)$, and the image curve $\beta(t)\coloneqq(\psi\circ\alpha)(t)=(\sqrt{a}x(t),\sqrt{b}y(t))$. Applying $G_{\mathcal{H}^{2}}(\beta(t))$ to $\dot{\beta}(t)=\mathrm{d}\psi_{\alpha(t)}(\dot{\alpha}(t))$ in $\mathcal{H}^{2}$ gives
@@
Thus, $\left\langle\dot{\alpha}(t),\dot{\alpha}(t)\right\rangle_{G_{\mathcal{M}}(\alpha(t))}=b\left\langle\dot{\beta}(t),\dot{\beta}(t)\right\rangle_{G_{\mathcal{H}^{2}}(\beta(t))}$, that is, $\left\|\dot{\alpha}(t)\right\|_{G_{\mathcal{M}}(\alpha(t))}=\sqrt{b}\left\|\dot{\beta}(t)\right\|_{G_{\mathcal{H}^{2}}(\beta(t))}$. This implies that a curve $\alpha(t)$ connecting two points in $\mathcal{M}$ is a geodesic (minimises the length) if, and only if, its image $\beta(t)$ is a geodesic in $\mathcal{H}^{2}$. Taking (1.7) and (1.8) into account concludes the proof. ∎
@@
that is, $\left\langle(u_{1},\dots,u_{n}),(v_{1},\dots,v_{n})\right\rangle_{G_{\mathcal{H}^{n}}(x_{1},\dots,x_{n})}=\frac{1}{x_{n}^{2}}\sum_{i=1}^{n}u_{i}v_{i}$. This manifold has constant negative curvature. The geodesics in this manifold are vertical half-lines and vertical half-circles centred at the hyperplane $x_{n}=0$, and the geodesic distance between two points in $\mathcal{H}^{n}$ is given by
@@
Restricted to points in the $(n-1)$-dimensional unit half-sphere in $\mathcal{H}^{n}$, that is, $S_{1}^{n-1}\coloneqq\left\{(x_{1},\dots,x_{n})\in\mathcal{H}^{n}\;\middle|\;\sum_{i=1}^{n}x_{i}^{2}=1\right\}$, the expression for the distance becomes
$\widetilde{d}_{S^{n-1}_{1}}\big{(}(x_{1},\dots,x_{n}),(y_{1},\dots,y_{n})\big{)}=\operatorname{arccosh}\left(\frac{1-\sum_{i=1}^{n-1}x_{i}y_{i}}{x_{n}y_{n}}\right).$ (2.5)
From (2.4), we immediately see that the distance $d_{\mathcal{H}^{n}}$ is invariant by dilation or contraction, that is, denoting $\boldsymbol{x}\coloneqq(x_{1},\dots,x_{n})\in\mathcal{H}^{n}$ and $\boldsymbol{y}\coloneqq(y_{1},\dots,y_{n})\in\mathcal{H}^{n}$, we have $d_{\mathcal{H}^{n}}(\boldsymbol{x},\boldsymbol{y})=d_{\mathcal{H}^{n}}(\lambda\boldsymbol{x},\lambda\boldsymbol{y})$, for $\lambda\neq 0$. For points $\boldsymbol{x}$ and $\boldsymbol{y}$ in the half-sphere $S_{r}^{n-1}\coloneqq\left\{(x_{1},\dots,x_{n})\in\mathcal{H}^{n}\;\middle|\;\sum_{i=1}^{n}x_{i}^{2}=r^{2}\right\}$ of radius $r$, we can see that the distance between them, restricted to the sphere $S_{r}^{n-1}$, is the distance given by (2.5) between their projections onto the radius-$1$ sphere, that is,
$\widetilde{d}_{S_{r}^{n-1}}\left(\boldsymbol{x},\boldsymbol{y}\right)=\widetilde{d}_{S_{1}^{n-1}}\left(\frac{\boldsymbol{x}}{r},\frac{\boldsymbol{y}}{r}\right).$ (2.6)